The average body temperature of a healthy person is 37.0°C. Boltzmann's constant is 1.38 × 10−23 J/K. If a fever increases the temperature to 38.0°C, by what percentage does the average kinetic energy of the molecules increase?

Answer must be in percentage.
Answer #1

AnswerGiven initial temperatuu- 37°C . = 273+37= 310k Final temperature = 38°C 273+ 38 = 311k Boltz man Constanta ko = 1.38 X 16 23